OpenStudy (anonymous):

This is kind of a long one. A perpendicular cut through a water pipe reveals a circular cross section. The diameter of a 3/4 inch pipe is 50% greater than the diameter of a 1/2 inch pipe. The cross-sectional area of the larger pipe is what percent greater than the cross sectional area of the smaller pipe?

OpenStudy (tkhunny):

Diameter of 3/4" Pipe = 1.50 * Diameter of 1/2" Pipe Radius of 3/4" Pipe = 1.50 * Radius of 1/2" Pipe \(Area = \pi r^{2}\)